OpenStudy (anonymous):

Connor went to New York for vacation. He spent 3 nights at a hotel and rented a car for 4 days. Jillian stayed at the same hotel, but spent 4 nights and rented a car for 5 days from the same company. If Connor paid $675 and Jillian paid $875, how much did one night at the hotel cost?

OpenStudy (campbell_st):

well this is a pair of linear equations that can be solved simultaneously Connor : 3h + 4c = 675 (1) Jillian : 4h + 5c = 875 (2) multiply equation (1) by 5 and (2) by 4 then subtract the equations. this will eliminate c 15h + 20c = 3375 - 16h + 20c = 3500 ------------------------ -h = - 125 I'll let you finish the solution. hope it helps

OpenStudy (anonymous):

125?